# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio

### v1

@@ Abstract
Solving large-scale, continuous-time portfolio optimization problems involving numerous assets and state-dependent dynamics has long been challenged by the curse of dimensionality. Traditional dynamic programming and PDE-based methods, while rigorous, typically become computationally intractable beyond a small number of state variables (often limited to  3-6 in prior numerical studies). To overcome this critical barrier, we introduce the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework. PG-DPO leverages Pontryagin’s Maximum Principle to directly guide neural network policies via backpropagation-through-time, naturally incorporating exogenous state processes without requiring dense state grids. Crucially, our computationally efficient “Two-Stage” variant exploits rapidly stabilizing costate estimates derived from BPTT, converting them into near-optimal closed-form Pontryagin controls after only a short warm-up, significantly reducing training overhead. This enables a breakthrough in scalability: numerical experiments demonstrate that PG-DPO successfully tackles problems with dimensions previously considered far out of reach, optimizing portfolios with up to 50 assets and 10 state variables. The framework delivers near-optimal policies, offering a practical and powerful alternative for high-dimensional continuous-time portfolio choice.
@@ 1 Introduction
A promising alternative is to employ Pontryagin’s Maximum Principle (PMP), an indirect method that bypasses the explicit computation of value functions (Pontryagin, , 2014; Yong & Zhou, , 2012). Rather than storing or approximating a high-dimensional value grid, the PMP characterizes the optimal control in terms of forward–backward differential equations for the state and its costate (adjoint). Notably, adding one more state dimension does not produce an exponential explosion in computational requirements, in contrast to DP (Bellman, , 1966; Campbell et al., , 2003). This makes Pontryagin-based methods appealing for large-scale portfolio optimization problems involving many assets and exogenous state processes (Karatzas et al., , 1987).
@@
Recent advances in deep learning further enhance the scalability of methods for solving high-dimensional stochastic control problems (e.g., Han et al., , 2018; Raissi et al., , 2019, for general PDE-based approaches). Concurrently, deep learning techniques are being applied directly to financial dynamic programming. Some approaches tackle infinite horizon problems via the HJB equation (Duarte et al., , 2024), while others aim to directly solve the associated nonlinear PDEs using methods like deep branching algorithms (Nguwi et al., , 2024). However, applications such as the Merton HJB have revealed limitations in accuracy or efficiency for certain PDE solvers like Deep BSDE variants (Nguwi et al., , 2024). Our approach, distinct from these direct HJB/PDE solvers, focuses on using Pontryagin’s principle, readily handling finite horizons. One notable strategy related to the use of adjoint methods in optimal control is to directly parameterize the policy (i.e., consumption and investment controls) with neural networks, then apply backpropagation-through-time (BPTT) to enforce Pontryagin’s optimality conditions (Huh, , 2024). This strategy, known as “Pontryagin-Guided Direct Policy Optimization (PG-DPO),” circumvents the need to approximate value functions or PDE solutions. Instead, it uses simulation-based gradients that automatically incorporate the adjoint processes, drastically reducing computational overhead in high-dimensional settings (Kushner & Yin, , 2003; Borkar & Borkar, , 2008).
@@
The purpose of this paper is to develop and implement this PG-DPO framework for the general multi-asset, state-dependent continuous-time portfolio problem. We consider an investor who must decide on a continuous consumption rate and an allocation to $n$ risky assets with state-dependent drift and volatility. By parameterizing the policy with neural networks, we can capture potentially complex relationships dependent on both wealth and the exogenous state. We show how a single forward pass in simulation, coupled with a backward pass via automatic differentiation (BPTT), yields unbiased estimates of the costate processes that PMP requires, guiding the policy towards optimality without the drawbacks of DP.
@@
Furthermore, we introduce a “Two-Stage” PG-DPO variant that exploits the observation that costate estimates often stabilize much faster than the full policy network. Once the costates are sufficiently stable (after a short warm-up), one can bypass the potentially suboptimal neural policy outputs and directly substitute the costate estimates into the PMP first-order conditions. This recovers a near-optimal control analytically at every time-state point with minimal additional training, significantly reducing computational burden while maintaining high accuracy. This efficiency enables the solution of problems with dozens of assets and state variables (e.g., n=50, k=10 in our experiments), a scale considerably larger than those typically addressed by the numerical methods cited previously.
@@
  • We develop and implement the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework, leveraging PMP and deep learning, for solving general multi-asset, multi-factor continuous-time portfolio optimization problems with state-dependent dynamics and finite horizons.
  • We introduce a computationally efficient “Two-Stage” PG-DPO variant that utilizes rapidly stabilized costate estimates from BPTT directly within the PMP first-order conditions, significantly reducing training overhead while achieving high accuracy.
  • We demonstrate through numerical experiments that our framework, particularly the Two-Stage variant, successfully scales to high-dimensional problems involving dozens of assets and state variables (up to n=50, k=10). This substantially exceeds the dimensionality manageable by traditional DP methods and tackled in many prior numerical studies focusing on problems with typically fewer than  5-6 state variables, thus effectively mitigating the curse of dimensionality.
In the remainder of this paper, we first present the multi-asset Merton problem with exogenous states in Section 2. We focus on the affine multi-factor OU model, detailed in Section 2.3.2, for developing the methodology and conducting numerical experiments due to the availability of a benchmark solution. Section 3 introduces our Pontryagin-Guided Direct Policy Optimization (PG-DPO) algorithm and its Two-Stage extension. Numerical experiments presented in Section 4 confirm the framework’s scalability and near-optimal performance for the affine model. Finally, Section 5 concludes and discusses potential future directions, including applications to non-affine models or problems with portfolio constraints.
@@ 2 Multi-Asset Continuous-Time Portfolio Problem with State Variables
In this section, we present a multi-asset extension of Merton’s continuous-time portfolio problem (Merton, , 1971), allowing the investor to allocate wealth across multiple (potentially correlated) risky assets and a risk-free asset, while consuming continuously. Each risky asset features its own drift and volatility, leading to an $n$-dimensional portfolio weight vector $\bm{\pi}_{t}$. Furthermore, an exogenous state process $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ influences market parameters such as returns and interest rates, adding additional complexity to the investor’s decision.
@@ 2.1 Problem Formulation with State Variables and Neural Network Controls
Let $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ be a vector of state variables that affect the distribution of asset returns. For instance, $\mathbf{Y}_{t}$ can include interest rates, macroeconomic indices, volatility factors, and so forth. We assume each component $Y_{i,t}$ follows an Ito diffusion
@@
We consider a financial market with $n$ risky assets and one risk-free asset. For $i=1,\dots,n$, the price $S_{i,t}$ of the $i$-th risky asset follows
@@
Here, $\mu_{i}(t,\mathbf{Y}_{t})$ is the instantaneous expected return and $\sigma_{i}(t,\mathbf{Y}_{t})$ is the instantaneous volatility of asset $i$. The $n$ scalar Brownian motions $W^{X}_{i,t}$ may be correlated amongst themselves ($\operatorname{Cov}(dW^{X}_{i,t},\,dW^{X}_{j,t})=\psi_{ij}\,dt$) and with the state variable processes ($\operatorname{Cov}(dW^{X}_{i,t},\,dW^{Y}_{j,t})=\rho_{ij}\,dt$). In vector notation, let $\bm{\mu}(t,\mathbf{Y}_{t})$ be the $n\times 1$ vector of expected returns, $\bm{\sigma}(t,\mathbf{Y}_{t})$ be the $n\times n$ diagonal matrix of volatilities, and $\mathbf{W}^{X}_{t}$ be the $n\times 1$ vector of Brownian motions. The risky asset dynamics can be written compactly as
$d\mathbf{S}_{t}=\bm{\mu}(t,\mathbf{Y}_{t})\,\mathbf{S}_{t}\,dt+\bm{\sigma}(t,\mathbf{Y}_{t})\,\mathbf{S}_{t}\,d\mathbf{W}^{X}_{t}.$
@@
Let $X_{t}$ denote the investor’s total wealth. At each time $t$, the investor chooses a consumption rate $C_{t}\geq 0$ and allocates fractions of wealth to the $n$ risky assets, represented by the vector $\bm{\pi}_{t}=(\pi_{1,t},\,\dots,\,\pi_{n,t})^{\top}\in\mathbb{R}^{n}.$ The remaining fraction $1-\bm{\pi}_{t}^{\top}\mathbf{1}$ is allocated to the risk-free asset, which earns interest at rate $r\bigl{(}t,\mathbf{Y}_{t}\bigr{)}$. The wealth process then evolves according to the stochastic differential equation (SDE):
$dX_{t}\;=\;\Bigl{[}\,X_{t}\,\Bigl{(}\bigl{(}1-\bm{\pi}_{t}^{\top}\mathbf{1}\bigr{)}\,r\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\;+\;\bm{\pi}_{t}^{\top}\,\bm{\mu}(t,\mathbf{Y}_{t})\Bigr{)}\;-\;C_{t}\Bigr{]}\,dt\;+\;X_{t}\,\bm{\pi}_{t}^{\top}\,\bm{\sigma}(t,\mathbf{Y}_{t})\,d\mathbf{W}^{X}_{t}.$
We require the controls $(\bm{\pi}_{t},C_{t})$ to be admissible, ensuring that the wealth remains strictly positive, $X_{t}>0$, for all $t\in[0,T]$. This condition is essential for the utility function to be well-defined and is typically maintained in simulations (e.g., by clamping wealth at a small positive threshold, see Section 4.1.2).
For clarity on the noise structure, we define the combined $(n+k)$-dimensional Brownian vector $\mathbf{W}_{t}=(\mathbf{W}^{X\top}_{t},\mathbf{W}^{Y\top}_{t})^{\top}$. The covariance structure of its increments $d\mathbf{W}_{t}$ is given by the block matrix:
$\operatorname{Cov}\bigl{(}d\mathbf{W}_{t}\bigr{)}\;=\;\begin{pmatrix}\Psi&\rho\\ \rho^{\top}&\Phi\end{pmatrix}\,dt,$
where $\Psi\in\mathbb{R}^{n\times n}=\operatorname{Cov}(d\mathbf{W}^{X}_{t})/dt$, $\Phi\in\mathbb{R}^{k\times k}=\operatorname{Cov}(d\mathbf{W}^{Y}_{t})/dt$, and $\rho\in\mathbb{R}^{n\times k}=\operatorname{Cov}(d\mathbf{W}^{X}_{t},d\mathbf{W}^{Y}_{t})/dt$.
The investor’s objective is to choose admissible control processes $(\bm{\pi},C)$ to maximize the expected discounted utility functional $J$:
@@
subject to the wealth dynamics. Here $\delta>0$ is the subjective discount rate, and $\kappa\geq 0$ weights the utility of terminal wealth (bequest motive). The utility function $U:(0,\infty)\to\mathbb{R}$ represents the investor’s preferences and is assumed to be strictly increasing, strictly concave, and satisfy the Inada conditions. A widely used specification is the Constant Relative Risk Aversion (CRRA) form:
@@
with $\gamma>0$ denoting the coefficient of relative risk aversion.
In our numerical approach, we parameterize the control policies using neural networks:
$\bm{\pi}_{t}\;=\;\bm{\pi}_{\theta}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)},\quad C_{t}\;=\;C_{\phi}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)},$
where $\theta$ and $\phi$ represent the trainable parameters of the respective neural networks.
2.2 Pontryagin’s Maximum Principle with a State-Dependent Market and Costate for $\mathbf{Y}_{t}$
We now consider the continuous-time portfolio problem from Section 2.1 under a market whose parameters depend on the exogenous state process $\{\mathbf{Y}_{t}\}_{0\leq t\leq T}$. The Pontryagin Maximum Principle (PMP) provides necessary conditions for optimality via a system of forward-backward stochastic differential equations (FBSDEs).
The state of the system consists of the investor’s wealth $X_{t}$ and the exogenous state variables $\mathbf{Y}_{t}\in\mathbb{R}^{k}$. We introduce the costate processes associated with these states: $\lambda_{t}\approx\frac{\partial V}{\partial x}(t,X_{t},\mathbf{Y}_{t})$ (scalar costate for wealth) and $\bm{\eta}_{t}\approx\frac{\partial V}{\partial\mathbf{y}}(t,X_{t},\mathbf{Y}_{t})$ (a $k$-dimensional costate vector for the state variables), where $V$ is the value function of the problem. The dynamics of the state and costate processes are described by the following FBSDE system.
Under an optimal control policy $\bigl{(}\bm{\pi}_{t}^{*},\,C_{t}^{*}\bigr{)}$, the state processes $(X_{t}^{*},\mathbf{Y}_{t}^{*})$ evolve according to the forward SDEs:
$dX_{t}^{*}$ $=\;\Bigl{[}X_{t}^{*}\,\Bigl{(}\,\bigl{(}1-\bm{\pi}_{t}^{*\top}\mathbf{1}\bigr{)}\,r\bigl{(}t,\mathbf{Y}_{t}^{*}\bigr{)}\;+\;\bm{\pi}_{t}^{*\top}\,\bm{\mu}\bigl{(}t,\mathbf{Y}_{t}^{*}\bigr{)}\Bigr{)}\;-\;C_{t}^{*}\Bigr{]}\,dt$
$\quad+\;X_{t}^{*}\,\bm{\pi}_{t}^{*\top}\,\bm{\sigma}(t,\mathbf{Y}_{t}^{*})\,d\mathbf{W}_{t}^{X},$ (1)
$d\mathbf{Y}_{t}^{*}$ $=\;\bm{\mu}_{Y}\bigl{(}t,\mathbf{Y}_{t}^{*}\bigr{)}\,dt\;+\;\bm{\sigma}_{Y}\bigl{(}t,\mathbf{Y}_{t}^{*}\bigr{)}\,d\mathbf{W}_{t}^{Y},\quad X_{0}^{*}\;=\;x_{0}>0,\quad\mathbf{Y}_{0}^{*}\;=\;y_{0}\in\mathbb{R}^{k}.$
Here $\bm{\sigma}(t,\mathbf{Y}_{t}^{*})$ is the $n\times n$ (diagonal) matrix of volatilities, and $\bm{\sigma}_{Y}(t,\mathbf{Y}_{t}^{*})$ is the $k\times k$ diffusion matrix for $\mathbf{Y}_{t}^{*}$. The Brownian motions $\mathbf{W}_{t}^{X}\in\mathbb{R}^{n}$ and $\mathbf{W}_{t}^{Y}\in\mathbb{R}^{k}$ have the covariance structure $\operatorname{Cov}(d\mathbf{W}_{t})=\begin{pmatrix}\Psi&\rho\\ \rho^{\top}&\Phi\end{pmatrix}dt.$
Simultaneously, the costate processes $(\lambda_{t}^{*},\bm{\eta}_{t}^{*})$ satisfy the backward SDEs (BSDEs):
$-d\lambda_{t}^{*}$ $=\;\mathcal{H}_{x}\bigl{(}t,X_{t}^{*},\mathbf{Y}_{t}^{*},\bm{\pi}_{t}^{*},C_{t}^{*},\lambda_{t}^{*},\bm{\eta}_{t}^{*}\bigr{)}\,dt\;-\;\mathbf{Z}_{t}^{\lambda X}\,d\mathbf{W}_{t}^{X}\;-\;\mathbf{Z}_{t}^{\lambda Y}\,d\mathbf{W}_{t}^{Y},$ (2)
$-d\bm{\eta}_{t}^{*}$ $=\;\mathcal{H}_{\mathbf{y}}\bigl{(}t,X_{t}^{*},\mathbf{Y}_{t}^{*},\bm{\pi}_{t}^{*},C_{t}^{*},\lambda_{t}^{*},\bm{\eta}_{t}^{*}\bigr{)}\,dt\;-\;\mathbf{Z}_{t}^{\eta X}\,d\mathbf{W}_{t}^{X}\;-\;\mathbf{Z}_{t}^{\eta Y}\,d\mathbf{W}_{t}^{Y}.$ (3)
Here, $\mathcal{H}_{x}$ and $\mathcal{H}_{\mathbf{y}}$ denote the partial derivatives of the Hamiltonian function $\mathcal{H}$ with respect to $x$ and $\mathbf{y}$, evaluated at the optimal state and controls. The terms $\mathbf{Z}_{t}^{\lambda X},\mathbf{Z}_{t}^{\lambda Y},\mathbf{Z}_{t}^{\eta X},\mathbf{Z}_{t}^{\eta Y}$ are adapted processes representing the diffusion coefficients of the costates, related to second derivatives of the value function and the system’s volatility. The terminal conditions are $\lambda_{T}^{*}=\kappa\,e^{-\delta T}\,U^{\prime}\bigl{(}X_{T}^{*}\bigr{)}$ and $\bm{\eta}_{T}^{*}=\mathbf{0}.$
The Hamiltonian $\mathcal{H}$ associated with this control problem, whose partial derivatives appear in the BSDE drift terms, is given by:
$\begin{split}\mathcal{H}(t,x,\mathbf{y},\bm{\pi},C,\lambda,\bm{\eta})=&e^{-\delta t}\,U(C)+\lambda\Bigl{[}x\bigl{(}(1-\bm{\pi}^{\top}\mathbf{1})r(t,\mathbf{y})+\bm{\pi}^{\top}\bm{\mu}(t,\mathbf{y})\bigr{)}-C\Bigr{]}\\ &+\bm{\eta}^{\top}\bm{\mu}_{Y}(t,\mathbf{y}).\end{split}$
Optimality requires that the controls $(\bm{\pi}_{t}^{*},C_{t}^{*})$ maximize the Hamiltonian at each time $t$. The first-order condition (FOC) with respect to consumption $C_{t}$ yields $e^{-\delta t}\,U^{\prime}(C_{t}^{*})-\lambda_{t}^{*}=0$, implying $C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*}).$ The FOC with respect to the portfolio weights $\bm{\pi}_{t}$ leads to a condition balancing expected excess returns against risk, involving the costate $\lambda_{t}^{*}$ and its sensitivity $\mathbf{Z}_{t}^{\lambda X}$. A standard form for this condition is:

### v2

@@ Abstract
Solving large-scale, continuous-time portfolio optimization problems involving numerous assets and state-dependent dynamics has long been challenged by the curse of dimensionality. Traditional dynamic programming and PDE-based methods, while rigorous, typically become computationally intractable beyond few state variables ($\sim$3-6 limit in prior studies). To overcome this critical barrier, we introduce the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework. PG-DPO leverages Pontryagin’s Maximum Principle (PMP) and backpropagation-through-time (BPTT) to guide neural network policies, handling exogenous states without dense grids. This PMP-guided approach holds potential for a broad class of sufficiently regular continuous-time control problems. Crucially, our computationally efficient “Two-Stage” variant exploits rapidly stabilizing BPTT costate estimates, converting them into near-optimal Pontryagin controls after only a short warm-up, significantly reducing training overhead. This enables a breakthrough in scalability: numerical experiments show PG-DPO successfully tackles problems with dimensions previously considered far out of reach (up to 50 assets and 10 state variables). The framework delivers near-optimal policies, offering a practical and powerful alternative for high-dimensional continuous-time portfolio choice.
@@ 1 Introduction
A promising alternative is to employ Pontryagin’s Maximum Principle (PMP), an indirect method that bypasses the explicit computation of value functions (Pontryagin, , 2014; Yong & Zhou, , 2012) Rather than storing or approximating a high-dimensional value grid, the PMP characterizes the optimal control in terms of forward–backward differential equations for the state and its costate (adjoint). Notably, adding one more state dimension does not produce an exponential explosion in computational requirements, in contrast to DP (Bellman, , 1966; Campbell et al., , 2003). This makes Pontryagin-based methods appealing for large-scale portfolio optimization problems involving many assets and exogenous state processes (Karatzas et al., , 1987).
@@
Recent advances in deep learning further enhance the scalability of methods for solving high-dimensional stochastic control problems (e.g., Han et al., , 2018; Raissi et al., , 2019, for general PDE-based approaches). Concurrently, deep learning techniques are being applied directly to financial dynamic programming. Some approaches tackle infinite horizon problems via the HJB equation (Duarte et al., , 2024), while others aim to directly solve the associated nonlinear PDEs using methods like deep branching algorithms (Nguwi et al., , 2024). However, applications such as the Merton HJB have revealed limitations in accuracy or efficiency for certain PDE solvers like Deep BSDE variants (Nguwi et al., , 2024). Our PMP-based PG-DPO framework offers a distinct path, naturally handling finite horizons and focusing on the optimality conditions derived from PMP, which are applicable to a wide range of stochastic control problems possessing sufficient regularity. One notable strategy related to the use of adjoint methods in optimal control is to directly parameterize the policy (i.e., consumption and investment controls) with neural networks, then apply backpropagation-through-time (BPTT) to enforce Pontryagin’s optimality conditions (Huh, , 2024). This strategy, known as “Pontryagin-Guided Direct Policy Optimization (PG-DPO),” circumvents the need to approximate value functions or PDE solutions. Instead, it uses simulation-based gradients that automatically incorporate the adjoint processes, drastically reducing computational overhead in high-dimensional settings (Kushner & Yin, , 2003; Borkar & Borkar, , 2008).
@@
The purpose of this paper is to develop and implement this PG-DPO framework for the general multi-asset, state-dependent continuous-time portfolio problem. We consider an investor who must decide on a continuous consumption rate and an allocation to $n$ risky assets with state-dependent drift and volatility. By parameterizing the policy with neural networks, we can capture potentially complex relationships dependent on both wealth and the exogenous state. We show how a single forward pass in simulation, coupled with a backward pass via automatic differentiation (BPTT), yields unbiased estimates of the costate processes that PMP requires, guiding the policy towards optimality without the drawbacks of DP. This approach fundamentally relies on the existence and accurate estimation of the costate processes dictated by PMP, applicable whenever the control problem satisfies appropriate regularity conditions allowing for a well-defined Hamiltonian system.
@@
Furthermore, we introduce a “Two-Stage” PG-DPO variant that exploits the observation that costate estimates often stabilize much faster than the full policy network. Once the costates are sufficiently stable (after a short warm-up), one can bypass the potentially suboptimal neural policy outputs and directly substitute the costate estimates into the PMP first-order conditions. This recovers a near-optimal control analytically at every time-state point with minimal additional training, significantly reducing computational burden while maintaining high accuracy. This efficiency enables the solution of problems with dozens of assets and state variables (e.g., n=50, k=10 in our experiments), a scale considerably larger than those typically addressed by the numerical methods cited previously. This demonstrated scalability on a benchmarkable problem highlights the potential of leveraging PMP structure via BPTT estimates for tackling other complex, high-dimensional problems that may exhibit similar amenability to costate estimation under suitable regularity conditions.
@@
  • We introduce Pontryagin-Guided Direct Policy Optimization (PG-DPO), featuring a highly efficient “Two-Stage” variant that uses PMP structure and BPTT-derived costates to solve complex continuous-time control problems.
  • We demonstrate through numerical experiments the framework’s ability to break the dimensional barrier, successfully tackling portfolio optimization problems with dozens of assets and state factors (up to n=50, k=10).
In the remainder of this paper, we first present the multi-asset Merton problem with exogenous states in Section 2. Section 3 introduces our Pontryagin-Guided Direct Policy Optimization (PG-DPO) algorithm and its Two-Stage extension. Numerical experiments presented in Section 4 confirm the framework’s scalability and near-optimal performance within this testbed, illustrating its potential for high-dimensional applications. Finally, Section 5 concludes and discusses potential future directions.
@@ 2 Multi-Asset Continuous-Time Portfolio Problem with State Variables
In this section, we present a multi-asset extension of Merton’s continuous-time portfolio problem (Merton, , 1971; Kim & Omberg, , 1996; Liu, , 2007), allowing the investor to allocate wealth across multiple (potentially correlated) risky assets and a risk-free asset, while consuming continuously. Each risky asset features its own drift and volatility, leading to an $n$-dimensional portfolio weight vector $\bm{\pi}_{t}$. Furthermore, an exogenous state process $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ influences market parameters such as returns and interest rates, adding additional complexity to the investor’s decision.
@@ 2.1 Problem Formulation with State Variables and Neural Network Controls
Let $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ be a vector of economic state variables that influence the distribution of asset returns. For instance, $\mathbf{Y}_{t}$ could include interest rates, macroeconomic indices, volatility factors, etc. We assume each component $Y_{i,t}$ follows an Ito diffusion:
@@
We consider a financial market with $n$ risky assets and one risk-free asset. The risk-free asset accumulates wealth at an instantaneous rate $r(t,\mathbf{Y}_{t})$, which may depend on time and the current state. For $i=1,\dots,n$, the price $S_{i,t}$ of the $i$-th risky asset follows
@@
Here, $\mu_{i}(t,\mathbf{Y}_{t})$ is the instantaneous expected return and $\sigma_{i}(t,\mathbf{Y}_{t})$ is the instantaneous volatility of asset $i$. The $n$ scalar Brownian motions $W^{X}_{i,t}$ may be correlated amongst themselves ($\operatorname{Cov}(dW^{X}_{i,t},\,dW^{X}_{j,t})=\psi_{ij}\,dt$) and with the state variable processes ($\operatorname{Cov}(dW^{X}_{i,t},\,dW^{Y}_{j,t})=\rho_{ij}\,dt$). In vector notation, let $\bm{\mu}(t,\mathbf{Y}_{t})$ be the $n\times 1$ vector of expected returns, $\bm{\sigma}(t,\mathbf{Y}_{t})$ be the $n\times n$ diagonal matrix of volatilities, and $\mathbf{W}^{X}_{t}$ be the $n\times 1$ vector of Brownian motions for the risky assets. The risky asset dynamics can be written compactly as
$d\mathbf{S}_{t}=\operatorname{diag}(\mathbf{S}_{t})\,\bm{\mu}(t,\mathbf{Y}_{t})\,dt+\operatorname{diag}(\mathbf{S}_{t})\,\bm{\sigma}(t,\mathbf{Y}_{t})\,d\mathbf{W}^{X}_{t}.$
Throughout, we work on a complete probability space $\bigl{(}\Omega,\mathcal{F},\mathbb{P}\bigr{)}$ that carries the $(n+k)$-dimensional Brownian motion $\mathbf{W}_{t}$ and its natural filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$, augmented to satisfy the usual conditions (right-continuity and completeness).
@@
The investor allocates wealth $X_{t}$ between the risk-free asset and the $n$ risky assets. Let $\bm{\pi}_{t}$ be the $n\times 1$ vector representing the fraction of wealth invested in each risky asset at time $t$. The investor also chooses a consumption rate $C_{t}$. The dynamics of the investor’s wealth $X_{t}$ are governed by the following stochastic differential equation (SDE):
$dX_{t}=\Bigl{[}\,X_{t}\Bigl{(}r(t,\mathbf{Y}_{t})+\bm{\pi}_{t}^{\top}\bigl{(}\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}\bigr{)}\Bigr{)}-C_{t}\Bigr{]}dt+X_{t}\,\bm{\pi}_{t}^{\top}\bm{\sigma}(t,\mathbf{Y}_{t})\,d\mathbf{W}^{X}_{t}.$ (1)
Here, $\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}$ represents the vector of instantaneous excess returns over the risk-free rate, and $\mathbf{1}$ is an $n\times 1$ vector of ones.
We impose regularity conditions on the market parameters.
Assumption 1 (Market-parameter regularity).
Fix a finite horizon $T>0$. The coefficients defining the state process $\mathbf{Y}_{t}$ and influencing the wealth dynamics $X_{t}$ via Eq. (1), namely $\bm{\mu}_{Y},\ \bm{\sigma}_{Y},\ \bm{\mu},\ \bm{\sigma},\ r$, satisfy the following conditions.
  1. (a)
    Measurability and continuity. Each map is jointly Borel-measurable in $(t,\mathbf{y})$ and continuous in $\mathbf{y}$ for every fixed $t\in[0,T]$.
  2. (b)
    Global Lipschitz continuity. There exists $K_{\mathrm{Lip}}>0$ such that for all $t\in[0,T]$ and $\mathbf{y},\mathbf{y}^{\prime}\in\mathbb{R}^{k}$,
    $\lVert\bm{\mu}_{Y}(t,\mathbf{y})-\bm{\mu}_{Y}(t,\mathbf{y}^{\prime})\rVert+\lVert\bm{\sigma}_{Y}(t,\mathbf{y})-\bm{\sigma}_{Y}(t,\mathbf{y}^{\prime})\rVert$
    $+\lVert\bm{\mu}(t,\mathbf{y})-\bm{\mu}(t,\mathbf{y}^{\prime})\rVert+\lVert\bm{\sigma}(t,\mathbf{y})-\bm{\sigma}(t,\mathbf{y}^{\prime})\rVert\leq K_{\mathrm{Lip}}\lVert\mathbf{y}-\mathbf{y}^{\prime}\rVert.$
  3. (c)
    Linear growth. There exists $K_{\mathrm{LG}}>0$ such that for all $(t,\mathbf{y})\in[0,T]\times\mathbb{R}^{k}$,
    $\lVert\bm{\mu}_{Y}(t,\mathbf{y})\rVert^{2}$ $+\lVert\bm{\sigma}_{Y}(t,\mathbf{y})\rVert^{2}+\lVert\bm{\mu}(t,\mathbf{y})\rVert^{2}+\lVert\bm{\sigma}(t,\mathbf{y})\rVert^{2}$
    $\leq K_{\mathrm{LG}}^{2}\bigl{(}1+\lVert\mathbf{y}\rVert^{2}\bigr{)},\qquad|r(t,\mathbf{y})|\leq K_{\mathrm{LG}}\bigl{(}1+\lVert\mathbf{y}\rVert\bigr{)}.$
Under Assumption 1, the SDE system for $(\mathbf{Y}_{t},X_{t})$ (where $X_{t}$ depends on the chosen controls $\bm{\pi}_{t},C_{t}$) admits a unique strong solution for sufficiently well-behaved controls; see Oksendal, (2013, Chs. 5-6) for details on existence and uniqueness theorems for SDEs.
We restrict our attention to control processes that are economically sensible and mathematically tractable.
Assumption 2 (Admissible controls).
A control pair $(\bm{\pi},C)$ is admissible if the following conditions hold:
  1. (a)
    Adaptedness. $(\bm{\pi}_{t},C_{t})_{0\leq t\leq T}$ is $\{\mathcal{F}_{t}\}$-progressively measurable, where $\{\mathcal{F}_{t}\}$ is the filtration generated by the underlying Brownian motions.
  2. (b)
    Integrability.
@@
  3. (c)
    Boundedness and proportional consumption. There exist constants $\Pi_{\max}>0$ and $\xi\in(0,1)$ such that
@@
  4. (d)
    Positive wealth. The corresponding wealth process $X_{t}$ solving (1) remains strictly positive almost surely: